# Llama-3.1-Sherkala-8B-Chat: An Open Large Language Model for Kazakh

### v1

@@ Appendix A Model Card
Model Dates Sherkala (8B) was trained between August 2024 and December 2024
@@
Intended Use Cases The Sherkala (8B) model is released with the aim to stimulate research and development in the Kazakh NLP community. It encourages researchers and hobbyists, especially those focusing on multi-lingual or non-English applications, to explore and to build upon the model. Feedback and collaboration opportunities are welcomed. The model is a pioneering addition to the Kazakh LLM ecosystem and has demonstrated exceptional Kazakh NLP capabilities compared to other open Kazakh or multilingual LLMs globally. Its applications span research advancements in Kazakh NLP, and the use of foundational models for fine-tuning.
Out-of-Scope Uses The Sherkala (8B) model is a powerful Kazakh and English language model with some capability in Russian, but it is important to recognize its limitations and the potential for misuse. Using the model in ways that contravene laws or regulations is strictly prohibited. This encompasses scenarios such as generating or endorsing hate speech, disseminating false information, engaging in illegal activities, managing sensitive data, attempting language generalization beyond Kazakh and English, and making critical decisions with high stakes. Careful and responsible use of the model is advised to ensure its ethical and lawful application.
@@
Overview Sherkala (8B) is trained from Llama-3.1 using 45.3 billion tokens, comprising 19.45 billion Kazakh tokens, 19.45 billion English tokens, and 6.4 billion Russian and Turkish tokens
@@
Table 13: Model card for Llama-3.1-Sherkala-8B-Chat.

### v2

@@ Appendix H Model Card
Model Dates Sherkala-Chat (8B) was trained between August 2024 and December 2024
@@
Intended Use Cases The Sherkala-Chat (8B) model is released with the aim to stimulate research and development in the Kazakh NLP community. It encourages researchers and hobbyists, especially those focusing on multi-lingual or non-English applications, to explore and to build upon the model. Feedback and collaboration opportunities are welcomed. The model is a pioneering addition to the Kazakh LLM ecosystem and has demonstrated exceptional Kazakh NLP capabilities compared to other open Kazakh or multilingual LLMs globally. Its applications span research advancements in Kazakh NLP, and the use of foundational models for fine-tuning.
Out-of-Scope Uses The Sherkala-Chat (8B) model is a powerful Kazakh and English language model with some capability in Russian, but it is important to recognize its limitations and the potential for misuse. Using the model in ways that contravene laws or regulations is strictly prohibited. This encompasses scenarios such as generating or endorsing hate speech, disseminating false information, engaging in illegal activities, managing sensitive data, attempting language generalization beyond Kazakh and English, and making critical decisions with high stakes. Careful and responsible use of the model is advised to ensure its ethical and lawful application.
@@
Overview Sherkala-Chat (8B) is trained from Llama-3.1 using 45.3 billion tokens, comprising 19.45 billion Kazakh tokens, 19.45 billion English tokens, and 6.4 billion Russian and Turkish tokens
@@
Table 12: Model card for Sherkala-Chat (8B).